(1 point) The matrix A = 1-6 1 8 k] 4 has two distinct real eigenvalues if and only if k > 24.5
Answer #1

* Answer: Given matrix A- f Characteristic (A-XI)=0 equation |-6-7 of K A is 1-0 8 4- * (-6-1) (4-X) -8k=0 *-24 +61-4x+x2-8k=